# Meta-World+: An Improved, Standardized, RL Benchmark

### v1

@@ Abstract
Meta-World is widely used for evaluating multi-task and meta-reinforcement learning agents, which are challenged to master diverse skills simultaneously. Since its introduction however, there have been numerous undocumented changes which inhibit a fair comparison of algorithms. This work strives to disambiguate these results from the literature, while also leveraging the past versions of Meta-World to provide insights into multi-task and meta-reinforcement learning benchmark design. Through this process we release a new open-source version of Meta-World***https://github.com/Farama-Foundation/Metaworld/ that has full reproducibility of past results, is more technically ergonomic, and gives users more control over the tasks that are included in a task set.
**footnotetext: These authors contributed equally to this work.
@@ 4 Meta-World
In this section we provide an overview of the Meta-World benchmark. In Appendix A we provide additional information on the different components of Meta-World. Meta-World contains 50 different robotic manipulation tasks for a single Sawyer Robot Arm***https://support.rethinkrobotics.com/support/solutions/articles/80000980284-arm-control-system to solve. Some tasks involve applying force to an object (i.e. door-close, drawer-close, and coffee-push), while some tasks require grasping and manipulating an object (i.e. pick-place, and assembly), and some combine these two applications of force (i.e. drawer-open, disassembly). In Appendix A we visualize the included tasks in Meta-World.
@@ 4.2 Reward Functions
The original publication of Meta-World (Yu et al., 2020b) created a set of dense reward functions. These rewards, which we will refer to as V1, were designed by creating the pick-place reward function, and then modifying the pick-place reward function to solve a new task. For the pick-place reward function, the rewards guide the agent to: reach towards an object, grasp the object, and move the object to a goal location. To create subsequent rewards this structure would be modified for each new task. For example, for the reach task, the reward function would be modified to only include the component of rewards for reaching towards the goal. The portion of the reward function for moving towards an object and gripping the object would be removed. However, at some point these V1 rewards were overwritten by a new set of dense rewards, which we will refer to as V2. These rewards were written with fuzzy constraints that allowed for rewards to be in the range $(0,10)$ and episodic returns across tasks are from a more narrow distribution. To highlight the differences in rewards, in Figure 1 we plot the per-timestep rewards for the pick-place task across the V1 and V2 rewards. For V1, the rewards start slightly negative and then reach a maximum reward for success of around 1200. For V2, the rewards start at zero and reach 10 when the task is solved. We include the full reasoning behind designing these reward functions in Appendix D.
@@ 5 Empirical Results
In order to compare the statistical results of each set of experiments we follow recommendations from Agarwal et al. (2021), where we report results over 10 random seeds and use interquartile mean (IQM) to compare results instead of simple averaging. For each graph we plot the IQM performance through training, while also including the 95% confidence interval. In multi-task learning, each method is evaluated for 50 evaluation episodes per task—once for each goal location. In meta-learning, the agent is evaluated after 10 adaptation episodes for 3 episodes for each goal location, per test task. Hyperparameters for each method are gathered from their respective publications. We implement each method into our open-source baselines code base using JAX (Bradbury et al. (2018)) ***https://github.com/rainx0r/metaworld-algorithms.
@@ A.4 Observation & Action Spaces
Both of the objects XYZ coordinates and quaternions can exist (i.e. in the coffee-push task), or if only one object exists the second object XYZ coordinates and quaternion are zeroed out. List A.4 only outlines an observation of size 18. The observation at time $t$ is stacked with the observation from time $t-1$, and the XYZ coordinates of the goal, to complete the 39-dimension observation vector. Meta-World is thus a goal-conditioned task. In multi-task reinforcement learning the goal is visible to the agent, while in meta-reinforcement learning the goal is zeroed out.
@@ A.5 Tasks & Task-Sets
{python}
import gymnasium as gym import metaworld # this registers the Meta-World environments with Gymnasium
# create a MT1 environment object envs = gym.make(’Meta-World/MT1’, env_name=’reach-v3’, seed=42)
# create a MT10 environment object envs = gym.make(’Meta-World/MT10’, vector_strategy=’sync’, seed=42)
# create a MT50 environment object envs = gym.make(’Meta-World/MT50’, vector_strategy=’sync’, seed=42)
# create a custom benchmark environment object envs = gym.make(’Meta-World/MT-custom’, env_names=[’env1-v3’, ’env2-v3’, …, ’envN-v3’], vector_strategy=’sync’, seed=42)
@@
{python}
import gymnasium as gym import metaworld # this registers the Meta-World environments with Gymnasium
# create ML1 environment objects train_envs = gym.make(’Meta-World/ML-1-train’, env_name=’reach-v3’, seed=42) test_envs = gym.make(’Meta-World/ML-1-test’, env_name=’reach-v3’, seed=42)
# create ML10 environment objects train_envs = gym.make(’Meta-World/ML10-train’, vector_strategy=’sync’, seed=42) test_envs = gym.make(’Meta-World/ML10-test’, vector_strategy=’sync’, seed=42)
# create a ML45 environment object train_envs = gym.make(’Meta-World/ML45-train’, vector_strategy=’sync’, seed=42) test_envs = gym.make(’Meta-World/ML45-test’, vector_strategy=’sync’, seed=42)
# create a custom ML benchmark environment object train_envs = gym.make(’Meta-World/ML-custom’, env_names=[’env1-v3’, ’env2-v3’, …, ’envN-v3’], vector_strategy=’sync’, seed=42) test_envs = gym.make(’Meta-World/ML-custom’, env_names=[’env1-v3’, ’env2-v3’, …, ’envN-v3’], vector_strategy=’sync’, seed=42)
@@
{python}
import gymnasium as gym import metaworld # this registers the Meta-World environments with Gymnasium from metaworld.evaluation import evaluation, metalearning_evaluation, Agent, MetaLearningAgent
# Multi-task # Explicitly subclassing is optional but recommended class MyAgent(Agent): def eval_action(self, observations): … # agent action selection logic
agent = MyAgent() envs = gym.make(’Meta-World/MT10’, vector_strategy=’sync’, seed=42) mean_success_rate, mean_returns, success_rate_per_task = evaluation(agent, envs)
# Meta-Learning # Explicitly subclassing is optional but recommended class MyMetaLearningAgent(MetaLearningAgent):
def eval_action(self, observations): # agent action selection logic …
def adapt_action(self, observations): # agent action selection logic … # this can also return values needed for adaptation like logprobs
def adapt(self, rollouts) -> None: # Adaptation logic here, this should mutate the agent object itself …
agent = MyMetaLearningAgent() test_envs = gym.make(’Meta-World/ML10-test’, vector_strategy=’sync’, seed=42) mean_success_rate, mean_returns, success_rate_per_task = metalearning_evaluation(agent, test_envs)
@@ Appendix C Replication
Table 2: Replication of selected MTRL results. For the PCGrad & SM results, these results use the V1 rewards. Both PaCo and MOORE leverage the V2 rewards. We note that our PCGrad results are not replicating the original PCGrad results, however it is inline with the values used in previous work [pmlr-v139-sodhani21a].
@@
Appendix D Reward Function Designs
D.1 Meta-World v1 Reward Functions
@@
D.2 Meta-World v2 Reward Functions
@@
D.3 Fuzzy-Logic Reward Function Paradigm

### v2

@@ Abstract
Meta-World is widely used for evaluating multi-task and meta-reinforcement learning agents, which are challenged to master diverse skills simultaneously. Since its introduction however, there have been numerous undocumented changes which inhibit a fair comparison of algorithms. This work strives to disambiguate these results from the literature, while also leveraging the past versions of Meta-World to provide insights into multi-task and meta-reinforcement learning benchmark design. Through this process we release a new open-source version of Meta-World [^1] https://github.com/Farama-Foundation/Metaworld/ that has full reproducibility of past results, is more technically ergonomic, and gives users more control over the tasks that are included in a task set.
@@ 4 Meta-World
In this section we provide an overview of the Meta-World benchmark. In Appendix A we provide additional information on the different components of Meta-World. Meta-World contains 50 different robotic manipulation tasks for a single Sawyer Robot Arm [^2] https://support.rethinkrobotics.com/support/solutions/articles/80000980284-arm-control-system to solve. Some tasks involve applying force to an object (i.e. door-close, drawer-close, and coffee-push), while some tasks require grasping and manipulating an object (i.e. pick-place, and assembly), and some combine these two applications of force (i.e. drawer-open, disassembly). In Appendix A we visualize the included tasks in Meta-World.
@@ 4.2 Reward Functions
The original publication of Meta-World (Yu et al., 2020b) created a set of dense reward functions. These rewards, which we will refer to as V1, were designed by creating the pick-place reward function, and then modifying the pick-place reward function to solve a new task. For the pick-place reward function, the rewards guide the agent to: reach towards an object, grasp the object, and move the object to a goal location. To create subsequent rewards this structure would be modified for each new task. For example, for the reach task, the reward function would be modified to only include the component of rewards for reaching towards the goal. The portion of the reward function for moving towards an object and gripping the object would be removed. However, at some point these V1 rewards were overwritten by a new set of dense rewards, which we will refer to as V2. These rewards were written with fuzzy constraints that allowed for rewards to be in the range $(0,10)$ and episodic returns across tasks are from a more narrow distribution. To highlight the differences in rewards, in Figure 1 we plot the per-timestep rewards for the pick-place task across the V1 and V2 rewards. For V1, the rewards start slightly negative and then reach a maximum reward for success of around 1200. For V2, the rewards start at zero and reach 10 when the task is solved. We include the full reasoning behind designing these reward functions in Appendix E.
@@ 5 Empirical Results
In order to compare the statistical results of each set of experiments we follow recommendations from Agarwal et al. (2021), where we report results over 10 random seeds and use interquartile mean (IQM) to compare results instead of simple averaging. For each graph we plot the IQM performance through training, while also including the 95% confidence interval. In multi-task learning, each method is evaluated for 50 evaluation episodes per task—once for each goal location. In meta-learning, the agent is evaluated after 10 adaptation episodes for 3 episodes for each goal location, per test task. Hyperparameters for each method are gathered from their respective publications. We implement each method into our open-source baselines code base using JAX (Bradbury et al. (2018))  [^3] https://github.com/rainx0r/metaworld-algorithms.
@@ 7 Conclusion
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract match the experimental results, and the contributions of this work are clearly outlined in the introduction.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: As we aim to provide guidelines for future benchmark design, we highlight some of the features that the Meta-World benchmark can add to aid research in additional areas of reinforcement learning research. These can be perceived as limitations of the current Meta-World benchmark.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: This work does not include theoretical results.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  15. Answer: [Yes]
  16. Justification: We will fully open source our new version of Meta-World, alongside the scripts we have used to generate the results of our experimental results.
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  19. Answer: [Yes]
  20. Justification: Anonymous URLs for code and environments are included in the submission, and will be publicly available upon acceptance. Exact scripts used to generate all results are available in the respective repo(s).
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  23. Answer: [Yes]
  24. Justification: Hyperparameter choices are described in the main text, and also available in our repo.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. Justification: We follow recommendations from Agarwal et al. [2021] and report results over 10 seeds using inter-quartile means and 95% confidence intervals in each plot.
  29. 8.
    Experiments compute resources
  30. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  31. Answer: [Yes]
  32. Justification: We discuss the computational resources needed for experimentation in the main text.
  33. 9.
    Code of ethics
  34. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  35. Answer: [Yes]
  36. Justification: We follow guidelines of Copyright and Fair Use.
  37. 10.
    Broader impacts
  38. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  39. Answer: [No]
  40. Justification: Our work focuses on the results of previous versions of the Meta-World benchmark, while providing a new version of Meta-World that is able to replicate past versions. While creating and maintaining a benchmark is an important task, we also note that benchmarks do cause research to focus on the improvement of benchmark scores rather than focusing on real-world deployments of reinforcement learning agents. This deployment of RL agents need to be generalizable, safe, and interpretable, none of which are the focus of our work.
  41. 11.
    Safeguards
  42. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  43. Answer: [N/A]
  44. Justification: We provide an updated version of a RL benchmark.
  45. 12.
    Licenses for existing assets
  46. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  47. Answer: [Yes]
  48. Justification: Licensing available in software repositories.
  49. 13.
    New assets
  50. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  51. Answer: [N/A]
  52. Justification:
  53. 14.
    Crowdsourcing and research with human subjects
  54. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  55. Answer: [N/A]
  56. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  57. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  58. Answer: [N/A]
  59. 16.
    Declaration of LLM usage
  60. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  61. Answer: [N/A]
@@ A.4 Observation & Action Spaces
Both of the objects XYZ coordinates and quaternions can exist (i.e. in the coffee-push task), or if only one object exists the second object XYZ coordinates and quaternion are zeroed out. List A.4 only outlines an observation of size 18. The observation at time $t$ is stacked with the observation from time t1t-1, and the XYZ coordinates of the goal, to complete the 39-dimension observation vector. Meta-World is thus a goal-conditioned task. In multi-task reinforcement learning the goal is visible to the agent, while in meta-reinforcement learning the goal is zeroed out.
@@ A.5 Tasks & Task-Sets
import metaworld
# this registers the Meta-World environments with Gymnasium
# create a MT1 environment object
envs = gym.make(’Meta-World/MT1’, env_name=’reach-v3’, seed=42)
# create a MT10 environment object
envs = gym.make(’Meta-World/MT10’, vector_strategy=’sync’, seed=42)
# create a MT50 environment object
envs = gym.make(’Meta-World/MT50’, vector_strategy=’sync’, seed=42)
# create a custom benchmark environment object
envs = gym.make(’Meta-World/MT-custom’, env_names=[’env1-v3’, ’env2-v3’, ..., ’envN-v3’], vector_strategy=’sync’, seed=42)
@@
import metaworld
# this registers the Meta-World environments with Gymnasium
# create ML1 environment objects
train_envs = gym.make(’Meta-World/ML-1-train’, env_name=’reach-v3’, seed=42)
test_envs = gym.make(’Meta-World/ML-1-test’, env_name=’reach-v3’, seed=42)
# create ML10 environment objects
train_envs = gym.make(’Meta-World/ML10-train’, vector_strategy=’sync’, seed=42)
test_envs = gym.make(’Meta-World/ML10-test’, vector_strategy=’sync’, seed=42)
# create a ML45 environment object
train_envs = gym.make(’Meta-World/ML45-train’, vector_strategy=’sync’, seed=42)
test_envs = gym.make(’Meta-World/ML45-test’, vector_strategy=’sync’, seed=42)
# create a custom ML benchmark environment object
train_envs = gym.make(’Meta-World/ML-custom’, env_names=[’env1-v3’, ’env2-v3’, ..., ’envN-v3’], vector_strategy=’sync’, seed=42)
test_envs = gym.make(’Meta-World/ML-custom’, env_names=[’env1-v3’, ’env2-v3’, ..., ’envN-v3’], vector_strategy=’sync’, seed=42)
@@
import metaworld
# this registers the Meta-World environments with Gymnasium
from metaworld.evaluation import evaluation, metalearning_evaluation, Agent, MetaLearningAgent
# Multi-task
# Explicitly subclassing is optional but recommended
class MyAgent(Agent):
def eval_action(self, observations):
... # agent action selection logic
agent = MyAgent()
envs = gym.make(’Meta-World/MT10’, vector_strategy=’sync’, seed=42)
mean_success_rate, mean_returns, success_rate_per_task = evaluation(agent, envs)
# Meta-Learning
# Explicitly subclassing is optional but recommended
class MyMetaLearningAgent(MetaLearningAgent):
def eval_action(self, observations):
# agent action selection logic
def adapt_action(self, observations):
# agent action selection logic
# this can also return values needed for adaptation like logprobs
def adapt(self, rollouts) -> None:
# Adaptation logic here, this should mutate the agent object itself
agent = MyMetaLearningAgent()
test_envs = gym.make(’Meta-World/ML10-test’, vector_strategy=’sync’, seed=42)
mean_success_rate, mean_returns, success_rate_per_task = metalearning_evaluation(agent, test_envs)
@@ Appendix C Replication
Table 2: Replication of selected MTRL results. For the PCGrad & SM results, these results use the V1 rewards. Both PaCo and MOORE leverage the V2 rewards. We note that our PCGrad results are not replicating the original PCGrad results, however it is inline with the values used in previous work [Sodhani et al., 2021].
@@
Appendix D Raw Results
Table 3: V1 MT10 Results
Algorithm Result
MTMHSAC 46.88 $\pm$ 22.69
Soft Modules 71.41 $\pm$ 2.08
MOORE 61.44 $\pm$ 7.10
PaCo 26.18 $\pm$ 7.54
PCGrad 70.49 $\pm$ 2.16
Table 4: V2 MT10 Results
Algorithm Result
MTMHSAC 77.26 $\pm$ 5.64
Soft Modules 84.88 $\pm$ 4.51
MOORE 83.42 $\pm$ 7.93
PaCo 72.26 $\pm$ 7.44
PCGrad 85.52 $\pm$ 1.72
Table 5: V1 MT50 Results
Algorithm Result
MTMHSAC 31.92 $\pm$ 2.67
Soft Modules 60.59 $\pm$ 3.86
MOORE 61.772 $\pm$ 2.66
PaCo 18.58 $\pm$ 15.38
PCGrad 45.84 $\pm$ 5.93
Table 6: V2 MT50 Results
Algorithm Result
MTMHSAC 54.90 $\pm$ 1.07
Soft Modules 65.78 $\pm$ 1.94
MOORE 71.99 $\pm$ 2.93
PaCo 58.88 $\pm$ 4.64
PCGrad 69.49 $\pm$ 2.67
Table 7: V1 ML10 Results
Algorithm Result
MAML V1 39.58 $\pm$ 16.11
RL2 V1 6.53 $\pm$ 7.74
Amago-2 V1 3 $\pm$ 4
Table 8: V2 ML10 Results
Algorithm Result
MAML V2 30.1 $\pm$ 11.45
MAML V2 (no baseline) 2.00 $\pm$ 3.46
RL2 V2 25.7 $\pm$ 8.92
Amago-2 V2 6 $\pm$ 4
Table 9: V1 ML45 Results
Algorithm Result
MAML V1 26.67 $\pm$ 7.67
RL2 V1 9.71 $\pm$ 12.31
Amago-2 V1 8 $\pm$ 4
Table 10: V2 ML45 Results
Algorithm Result
MAML V2 32.4 $\pm$ 14.47
RL2 V2 36.76 $\pm$ 13.90
Amago-2 V2 11 $\pm$ 3
Appendix E Reward Function Designs
E.1 Meta-World v1 Reward Functions
@@
E.2 Meta-World v2 Reward Functions
@@
E.3 Fuzzy-Logic Reward Function Paradigm